Exercise 9-13 (Part Level Submission) (Video)

Fultz Company has accumulated the following budget data for the year 2020.
1. Sales: 31,160 units, unit selling price $88.
2. Cost of one unit of finished goods: direct materials 1 pound at $6 per pound, direct labor 3 hours at $13 per hour, and manufacturing overhead $6 per direct labor hour.
3. Inventories (raw materials only): beginning, 10,280 pounds; ending, 15,490 pounds.
4. Selling and administrative expenses: $170,000; interest expense: $30,000.
5. Income taxes: 30% of income before income taxes.

Expand question part

(a)

Collapse question part

(b)

Prepare a budgeted multiple-step income statement for 2020.
FULTZ COMPANY
Budgeted Income Statement

For the Year Ending December 31, 2020December 31, 2020For the Quarter Ending December 31, 2020

Operating ExpensesNet Income / (Loss)Ending InventoryIncome from OperationsIncome Tax ExpensePurchasesSalesTotal Operating ExpensesSelling and Administrative ExpensesCost of Goods SoldGross ProfitIncome before Income TaxesBeginning Inventory

$

Total Operating ExpensesNet Income / (Loss)Ending InventorySalesBeginning InventoryOperating ExpensesCost of Goods SoldIncome Tax ExpenseIncome from OperationsGross ProfitIncome before Income TaxesPurchasesSelling and Administrative Expenses

Income from OperationsGross ProfitIncome before Income TaxesNet Income / (Loss)Beginning InventorySelling and Administrative ExpensesCost of Goods SoldEnding InventoryIncome Tax ExpenseOperating ExpensesPurchasesTotal Operating ExpensesSales

Total Operating ExpensesIncome before Income TaxesIncome from OperationsGross ProfitBeginning InventoryIncome Tax ExpenseCost of Goods SoldNet Income / (Loss)Operating ExpensesEnding InventoryPurchasesSalesSelling and Administrative Expenses

Total Operating ExpensesIncome before Income TaxesPurchasesBeginning InventorySalesSelling and Administrative ExpensesCost of Goods SoldEnding InventoryGross ProfitIncome from OperationsIncome Tax ExpenseNet Income / (Loss)Operating Expenses

PurchasesIncome before Income TaxesCost of Goods SoldEnding InventoryGross ProfitIncome from OperationsSelling and Administrative ExpensesSalesIncome Tax ExpenseNet Income / (Loss)Beginning InventoryInterest ExpenseTotal Operating Expenses

Income from OperationsIncome Tax ExpenseTotal Operating ExpensesSelling and Administrative ExpensesGross ProfitBeginning InventoryNet Income / (Loss)Cost of Goods SoldEnding InventoryOperating ExpensesIncome before Income TaxesPurchasesSales

Net Income / (Loss)Gross ProfitTotal Operating ExpensesIncome Tax ExpenseOperating ExpensesPurchasesCost of Goods SoldSalesIncome before Income TaxesSelling and Administrative ExpensesBeginning InventoryEnding InventoryIncome from Operations

Income from OperationsGross ProfitNet Income / (Loss)Beginning InventoryCost of Goods SoldEnding InventoryIncome Tax ExpenseIncome before Income TaxesPurchasesOperating ExpensesSalesSelling and Administrative ExpensesTotal Operating Expenses

$
Answer #1
FULTZ COMPANY
Budgeted Income Statement
For the Year Ending December 31, 2020
Sales 2742080 =31160*88
Cost of Goods Sold 1963080 =31160*63
Gross Profit 779000
Selling and Administrative Expenses 170000
Income from Operations 609000
Interest Expense 30000
Income before Income Taxes 579000
Income Tax Expense 173700
Net Income / (Loss) 405300
Workings:
Direct materials per unit 6 =1*6
Direct labor per unit 39 =3*13
Overhead per unit 18 =3*6
Cost of Goods Sold per unit 63